5. Determine the z-score value in each of the following scenarios:
a. What z-score value separates the top 8% of a normal distribution from the bottom
92%?
b. What z-score value separates the top 72% of a normal distribution from the bottom
28%?
c. What z-score value form the boundaries for the middle 58% of a normal
distribution?
d. What z-score value separates the middle 45% from the rest of the distribution?
1
Expert's answer
2019-05-15T13:40:53-0400

Consider normal distribution:


"p(x) = \frac{1}{\sqrt{2\pi\sigma^2}} \exp\bigg(-\frac{x^2}{2\sigma^2}\bigg)"


and


"F(x) = \int_{-\infty}^{x} p(t) \, dt"


and the probability to find x in (a,b) can be found as


"P(a<x<b) = F(b)-F(a)"

and

"P(x<b) = F(b)"

due to

"F(-\infty) = 0"


Then we need to find the solutions for

a) P(x<z) = F(z) = 92%

b) P(x<z) = F(z) = 28%

c) P(-z<x<z) = F(z) - F(-z) = 58%

d) P(-z<x<z) = F(z) - F(-z) = 45%

Here

"z = X \sigma"

where X is found numerically (in Mathematica) to be equal

a) X = 1.406

b) X = -0.583

c) X = 0.806

d) X = 0.598